OpenStudy (anonymous):

The solution to -45 = n over 5 is _____. -9 9 225 -225

OpenStudy (anonymous):

-45=n/5 so n=-45 multiply 5 so n=-225

OpenStudy (anonymous):

\[-45 = \frac{ n }{ n }\] do you mean this?

OpenStudy (anonymous):

The solution to -35x = -105 is -3. True False

OpenStudy (anonymous):

35 times 1 is 35 35 times 2 is 70 35 times 3 is 105

OpenStudy (anonymous):

The solution to w over -7 = -21 is 3. True False

OpenStudy (anonymous):

true

OpenStudy (anonymous):

false it will be 3 not -3 because -35x=-105 x=-105/-35=3(- on numerator and denominator cancels out)

OpenStudy (anonymous):